# The Complexity of Sequential Prediction in Dynamical Systems

### v1

@@ 3.1 Characterizations of Realizable Learnability
Starting with (a), let $t\in[T]$ such that $\texttt{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}\neq x_{t}$. We need to show that $\texttt{C}_{T-t}(V_{t+1},x_{t})<\texttt{C}_{T-t+1}(V_{t},x_{t-1})$. Suppose for the sake of contradiction that $\texttt{C}_{T-t}(V_{t+1},x_{t})\geq\texttt{C}_{T-t+1}(V_{t},x_{t-1}):=d$. Then, by the prediction rule, we have that $\texttt{C}_{T-t}(V^{\hat{x}_{t}}_{t},\hat{x}_{t})\geq\texttt{C}_{T-t}(V^{x_{t}}_{t},x_{t})\geq d$. By definition of the instance-dependent Evolution complexity, we are guaranteed the existence of a trajectory tree $\mathcal{T}_{x_{t}}$ of depth $T-t$, rooted at $x_{t}$, shattered by $V_{t}^{x_{t}}$ with branching factor at least $d$ and a tree $\mathcal{T}_{\hat{x}_{t}}$ of depth $T-t$, rooted at $\hat{x}_{t}$, shattered by $V_{t}^{\hat{x}_{t}}$ with branching factor at least $d$. Consider a binary tree $\mathcal{T}$ whose root node $v$ is labeled by $x_{t-1}$ and left and right subtrees are $\mathcal{T}_{x_{t}}$ and $\mathcal{T}_{\hat{x}_{t}}$ respectively. Then, observe that $\mathcal{T}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $d+1$ (because $x_{t}\neq\hat{x}_{t}$). This contradicts our assumption that $\texttt{C}_{T-t+1}(V_{t},x_{t-1})=d$. Thus, we must have $\texttt{C}_{T-t}(V_{t+1},x_{t})<\texttt{C}_{T-t+1}(V_{t},x_{t-1})$.
@@
Moving to (b), let $t\in[T]$ be such that $\hat{x}_{t}=x_{t}$. Then, we need to show that $d:=\texttt{C}_{T-t}(V_{t+1},x_{t})\leq\texttt{C}_{T-t+1}(V_{t},x_{t-1}).$ If $d=0$, the inequality holds trivially. Thus, assume $d>0$. Then, by definition, there exists a trajectory tree $\mathcal{T}$ of depth $T-t$, rooted at $x_{t}$, shattered by $V_{t+1}$ with branching factor at least $d$. Consider a binary tree $\tilde{\mathcal{T}}$ whose root node $v$ is labeled by $x_{t-1}$ and left and right subtrees are both $\mathcal{T}$. Then since $V_{t+1}=V_{t}^{x_{t}}\subseteq V_{t}$, we have that $\tilde{\mathcal{T}}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $d$. Thus, we must have $\texttt{C}_{T-t+1}(V_{t},x_{t-1})\geq d$.
@@
Finally for (c), let $t\in[T]$ be such that $\texttt{C}_{T-t+1}(V_{t},x_{t-1})=0$. Using the same logic as (b), $\texttt{C}_{T-t}(V_{t+1},x_{t})=0$. Thus, to prove that (2) holds, it suffices to show that $x_{t}=\hat{x}_{t}$. To do so, we will show that the projection size |$|\{f(x_{t-1}):f\in V_{t}\}|=1$. Thus, Algorithm 1 does not make a mistake. Suppose for the sake of contradiction that |$|\{f(x_{t-1}):f\in V_{t}\}|\geq 2$. Then, there exists two functions $f_{1},f_{2}\in V_{t}$ such that $f_{1}(x_{t-1})\neq f_{2}(x_{t-1})$. Consider a binary tree $\mathcal{T}$ of depth $T-t+1$ where the root node is labeled by $x_{t-1}$, and its left and right child by $f_{1}(x_{t-1})$ and $f_{2}(x_{t-1})$ respectively. Label all remaining internal nodes in the left subtree of the root node such that every path is shattered by $f_{1}$. Likewise, for the right subtree of the root node, label every internal node such that every path is shattered by $f_{2}$. Following this procedure, $\mathcal{T}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $1$ (because $f_{1}(x_{t-1})\neq f_{2}(x_{t-1})$). Thus, $\texttt{C}_{T-t+1}(V_{t},x_{t-1})\geq 1$, which is a contradiction. ∎
While Theorem 3.1 shows that the finiteness of $\texttt{E}_{\gamma}(\mathcal{F})$ for every $\gamma>0$ is sufficient for realizable learnability, it is not clear from the lowerbound that it is necessary. Theorem 3.2, whose proof is in Appendix C, resolves this gap by showing that indeed the finiteness of $\texttt{E}_{\gamma}(\mathcal{F})$ for every $\gamma>0$ is both necessary and sufficient for realizable learnability, thus providing a qualitative characterization.
Theorem 3.2 (Qualitative characterization of Realizable Learnability).
For every $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, the following statements are equivalent.
  • (i)
    $\mathcal{F}$ is learnable in the realizable setting.
  • (ii)
[…]
  • (iii)
    $\emph{{E}}_{\gamma}(\mathcal{F})$ is finite for every $\gamma>0$.
3.2 Minimax Rates in the Realizable Setting
While Theorems 3.1 and 3.2 provide a quantitative and qualitative characterization of realizable learnability, they do not shed light on how $\inf_{\mathcal{A}}\texttt{M}_{\mathcal{A}}(\mathcal{F},T)$ may depend on the time horizon $T$. In online classification with the 0-1 loss, the seminal work by Littlestone (1987) and Daniely et al. (2011) show that only two rates are possible: $\Theta(T)$ and $\Theta(1)$. That is, if a hypothesis class is online learnable in the realizable setting, then it is learnable with a constant mistake bound (i.e. the Littlestone dimension). Perhaps surprisingly, this is not the case for learning dynamical systems in a strong sense: every rate is possible.
Theorem 3.3.
For every $S\subset\mathbb{N}\cup\{0\}$, there exists $\mathcal{F}_{S}\subseteq\mathcal{\mathbbm{Z}}^{\mathbbm{Z}}$ such that $\emph{{C}}_{T}(\mathcal{F}_{S})=\sup_{n\in\mathbb{N}\cup\{0\}}|S\cap\{n,n+1,...,n+T-1\}|.$
Theorem 3.3, proved in Appendix D, along with Theorem 3.1 implies that any minimax rate in the realizable setting is possible. As an example, suppose we would like to achieve the rate $\Theta(T^{\alpha})$ for some $\alpha<1$. Then, picking $S=\{\lfloor t^{\frac{1}{\alpha}}\rfloor:t\in\mathbb{N}\cup\{0\}\}$ suffices since $\sup_{x\in\mathbb{N}\cup\{0\}}|S\cap\{n,n+1,...,n+T-1\}|=|\{\lfloor t^{\frac{1}{\alpha}}\rfloor:t\in\mathbb{N}\cup\{0\}\}\cap\{0,1,...,T-1\}|=\Theta(T^{\alpha})$. Likewise, one can get logarithmic rates by picking $S=\{2^{t}:t\in\mathbb{N}\cup\{0\}\}$ and constant rates by picking $S\subset\mathbb{N}\cup\{0\}$ such that $|S|<\infty$.
In light of Theorem 3.3 and the fact that only constant mistake bounds are possible for online multiclass classification, it is natural to ask when one can achieve constant mistake bounds for learning dynamical systems. To answer this question, we introduce a new combinatorial dimension termed the Branching dimension.
Definition 7 (Branching dimension).
The Branching dimension, denoted $\emph{{dim}}(\mathcal{F})$, is the smallest natural number $d\in\mathbb{N}$ such that for every shattered trajectory tree $\mathcal{T}$, we have $\emph{{B}}(\mathcal{T})\leq d$. If for every $d\in\mathbb{N}$, there exists a shattered trajectory tree $\mathcal{T}$ with $\emph{{B}}(\mathcal{T})>d$, we say $\emph{{dim}}(\mathcal{F})=\infty$.
Theorem 3.4, proved via non-constructive arguments in Appendix E, shows that $\inf_{\mathcal{A}}\texttt{M}_{\mathcal{A}}(T,\mathcal{F})=\Theta(1)$ if and and only if $\texttt{dim}(\mathcal{F})<\infty$.
Theorem 3.4 (Constant Minimax Expected Mistakes).
For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, we have
  • (i)
    $\inf_{\mathcal{A}}\emph{{M}}_{\mathcal{A}}(T,\mathcal{F})\leq\emph{{dim}}(\mathcal{F}).$
  • (ii)
    If $\emph{{dim}}(\mathcal{F})=\infty$, then $\inf_{\mathcal{A}}\emph{{M}}_{\mathcal{A}}(T,\mathcal{F})=\omega(1).$ [^1] Recall that $f(n)=\omega(1)$ if for all $k>0$, there exists $n_{k}$, such that for all $n>n_{k}$ we have $f(n)>k$.
3.3 Relations to PAC and Online Multiclass Classification
In this section, we study the relationship between realizable learnability of dynamical systems to realizable learnability in the well-known PAC and online classification settings. Our main result relates the evolution complexity and dimension to the DS and Littlestone dimensions.
Theorem 3.5 (Relations to the DS and Littlestone dimension).
The following statements are true.
  • (i)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\emph{{DS}}(\mathcal{F})=\infty$ but $\emph{{C}}_{T}(\mathcal{F})=\Theta(\log(T)).$
  • (ii)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\emph{{DS}}(\mathcal{F})=1$ but $\emph{{C}}_{T}(\mathcal{F})=T$.
  • (iii)
    For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$, we have that $\emph{{C}}_{T}(\mathcal{F})\leq\emph{{L}}(\mathcal{F})$.
  • (iv)
    There exists $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that $\emph{{L}}(\mathcal{F})=\infty$ but $\emph{{C}}_{T}(\mathcal{F})=1$.
The proof of Theorem 3.5 is in Appendix F. Parts (i) and (ii) show that the finiteness of the DSdim is neither necessary nor sufficient for learning dynamical systems in the realizable setting. On the other hand, parts (iii) and (iv) show that finite Ldim is sufficient but not necessary for learning dynamical systems in the realizable setting. Overall, learning dynamical systems is always easier than online multiclass classification, but can be both easier and harder than multiclass PAC classification. The proof of (i), (ii), and (iv) are combinatorial in nature, while the proof of (iii) involves reducing learning-to-predict in dynamical systems to online multiclass classification.
3.4 Examples
In this section, we establish the minimax rates for discrete linear systems and linear Boolean networks. Theorem 3.6 provides bounds on the Evolution complexity for discrete linear systems.
Theorem 3.6 (Discrete Linear Systems).
Let $\mathcal{X}=\mathbb{Z}^{n}$ and $r<n$. For =$\mathcal{F}=\{x\mapsto Wx\,:\,W\in\mathbb{Z}^{n\times n},\emph{\text{ rank}}(W)\leq r\}$ and $T>r$, we have $\emph{{C}}_{T}(\mathcal{F})=r+1.$
The proof of the upperbound in Theorem 3.6 uses the fact that given any initial state, the stream generated by any $f\in\mathcal{F}$ can have at most $r$ other states which are linearly independent of previous states. As for the lowerbound, we show how to construct a hard stream for every deterministic learner. Theorem 3.7, whose proof is in Appendix G.2, provides bounds on the Evolution complexity for two natural families of Boolean networks.
Theorem 3.7 (Linear Boolean Networks).
Let $\mathcal{X}=\{0,1\}^{n}$ and $T\geq n$.
  • (i)
    For =$\mathcal{F}=\{x\mapsto Wx\,(\emph{\text{mod }}2)\,:\,W\in\mathbb{Z}^{n\times n}\}$, we have $\emph{{C}}_{T}(\mathcal{F})=n$.
  • (ii)
    For =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{Wx>0\}\,:\,W\in\{0,1\}^{n\times n}\}$, we have $n\leq\emph{{C}}_{T}(\mathcal{F})\leq n^{2}$.
Thresholded Boolean networks have been used to model genetic regulatory dynamics (Mendoza and Alvarez-Buylla, 1998) and social networks (Kempe et al., 2003). Modulo Boolean networks have been studied by Chandrasekhar et al. (2023) in the context of stability. Similar to Theorem 3.6, the proof of (i) in Theorem 3.7 uses the fact that any realizable stream can have at most $n-1$ states (excluding the initial state) that are linearly independent of all preceeding states. To prove the upperbound in (ii), we bound the Ldim of $\mathcal{F}$. For the lowerbound, we construct a hard stream such that every deterministic learner makes at least $n$ mistakes. We conjecture that the upperbound in Theorem 3.7 (ii) can be reduced to $O(n)$, but we leave this as an interesting open question.
4 Agnostic Learnability under Markovian Regret
Theorem 4.1 (Minimax Expected Markovian Regret).
For any $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$,
$\max\left\{\frac{\emph{{L}}(\mathcal{F})}{18},\frac{\sqrt{T}}{16\sqrt{3}}\right\}\leq\inf_{\mathcal{A}}\,\emph{{R}}^{\emph{{MK}}}_{\mathcal{A}}(T,\mathcal{F})\leq\emph{{L}}(\mathcal{F})+\sqrt{T\,\emph{{L}}(\mathcal{F})\log{T}}.$
One such class exhibiting the separation is the thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\geq a\}\,:\,a\in(0,1)\}$ used in the proof of part (iv) in Theorem 3.5. Beyond the qualitative separation of realizable and agnostic learnability under Markovian regret, we also observe a quantitative separation in terms of possible minimax rates. Recall that Theorem 3.3 shows that every rate is possible in the realizable setting. However, Theorem 4.1 shows that only two types of rates are possible in the agnostic setting: $\Theta(T)$ whenever $\texttt{L}(\mathcal{F})=\infty$ and $\tilde{\Theta}(\sqrt{T})$ whenever $\texttt{L}(\mathcal{F})<\infty$. More precisely, when $\sqrt{T}\geq\texttt{L}(\mathcal{F})$, we have a lowerbound of $\Omega(\sqrt{T})$ and an upperbound of $O(\sqrt{T\texttt{L}(\mathcal{F})\log{T}})$ in Theorem 4.1. This raises the natural question of what the right minimax rate is. In Appendix I, we provide an evolution class and establish a lowerbound of $\Omega(\sqrt{T\texttt{L}(\mathcal{F})})$, showing that the upperbound is tight up to $\sqrt{\log{T}}$.
Our proof of the upperbound in Theorem 4.1 reduces learning dynamical systems to online multiclass classification and uses a result due to Hanneke et al. (2023). To prove the lowerbound $\frac{\texttt{L}(\mathcal{F})}{18}$, we construct a hard stream by carefully sampling a random path down a Littlestone tree of depth $\texttt{L}(\mathcal{F})$. To prove the lowerbound of $\frac{\sqrt{T}}{16\sqrt{3}}$, we construct a hard randomized stream using just two different evolution functions in $\mathcal{F}$. The full proof is in Appendix H.
5 Agnostic Learnability under Flow Regret
The necessity of the Littlestone dimension for agnostic learnability under Markovian regret is quite restrictive. One can verify that even for simple (but unnatural) classes like one dimensional thresholds =$\mathcal{F}=\{x\mapsto\mathbbm{1}\{x\leq a\}:a\in(0,1)]\}$, we have that $\texttt{L}(\mathcal{F})=\infty$ while $\texttt{C}_{T}(\mathcal{F})=1$. Indeed, the intuition in the lowerbound of Theorem 5.1 is that, in the agnostic setting, the adversary can force the learner to play an online multiclass classification game by “giving up” every other round. Since the Ldim is necessary for online multiclass classification, it is not surprising that it is also necessary for learning dynamical systems in the agnostic setting.
The main reason behind the equivalence of online multiclass classification and learning dynamical systems in the agnostic setting is the definition of Markovian regret. In particular, the evaluation of the “best-fixed evolution function in hindsight” does not penalize the evolution function beyond a one-step prediction error. From this perspective, perhaps a more natural definition of expected regret in the agnostic setting is to compare the prediction of the learner to the prediction of the best-fixed trajectory generated by functions in our evolution class. More rigorously, define
$\texttt{R}^{\texttt{FL}}_{\mathcal{A}}(T,\mathcal{F}):=\sup_{(x_{0},x_{1},\ldots,x_{T})}\Bigl{(}\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}\{\mathcal{A}(x_{<t})\neq x_{t}\}\right]-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\mathbbm{1}\{f^{t}(x_{0})\neq x_{t}\}\Bigl{)}.$ (3)
Definition 8 (Agnostic Learnability under Flow Regret).
[…]
Which evolution classes $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ are agnostic learnable under Flow regret? Is the finiteness of Ldim still necessary? Theorem 5.1, whose proof can be found in Appendix J, provides partial answers to these question by bounding the minimax expected Flow regret for classes where $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|$ is uniformly bounded.
Theorem 5.1 (Minimax Expected Flow Regret).
$\frac{\emph{{C}}_{T}(\mathcal{F})}{2}\leq\inf_{\mathcal{A}}\emph{{R}}^{\emph{{FL}}}_{\mathcal{A}}(T,\mathcal{F})\leq\emph{{C}}_{T}(\mathcal{F})+\sqrt{\emph{{C}}_{T}(\mathcal{F})\,T\ln\Bigl{(}\frac{T\,K_{\mathcal{F}}}{\emph{{C}}_{T}(\mathcal{F})}\Bigl{)}}.$
Note that the upperbound becomes vacuous when $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|=\infty$. We leave it as an open question to find the right characterization under Flow regret. Unlike agnostic learnability under Markovian regret, the finitness of the Ldim is not necessary for agnostic learnability under Flow regret. Indeed, while the class of one-dimensional thresholds is not agnostic learnable under Markovian regret, it is agnostic learnable under Flow regret. In fact, a separation between agnostic learnability for Markovian and Flow regret can be exhibited by any class $\mathcal{F}$ for which $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|<\infty$, $\texttt{L}(\mathcal{F})=\infty$, and $\texttt{C}_{T}(\mathcal{F})=o(T)$.
Theorem 5.1 shows that realizable and agnostic learnability under Flow regret are equivalent as long as $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|<\infty$. A natural question to ask is whether this is also true when $\sup_{x\in\mathcal{X}}|\mathcal{F}(x)|=\infty$. Unfortunately, Theorem 5.2 shows that when the projection sizes are not uniformly bounded, realizable and agnostic learnability under Flow regret are not equivalent.
There exists an ordered set $\mathcal{X}$ and $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that
  • (i)
    $\inf_{\mathcal{A}}\emph{{M}}_{\mathcal{A}}(T,\mathcal{F})\leq 3$.
  • (ii)
    $\inf_{\mathcal{A}}\emph{{R}}^{\emph{{FL}}}_{\mathcal{A}}(T,\mathcal{F})\geq\frac{T}{6}$.
To prove Theorem 5.2, we construct a class $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ such that on large subset of states in $\mathcal{X}^{\prime}\subset\mathcal{X}$, every function $f\in\mathcal{F}$ effectively reveals its identity. The full proof of Theorem 5.2 can be found in Appendix K.
6 Discussion and Future directions
Acknowledgements
We thank Chinmaya Kaushik for pointing out a technical result, which helped us to prove Theorem 3.6.
Appendix A Combinatorial dimensions
Definition 9 (DS dimension [Daniely and Shalev-Shwartz, 2014]).
We say that $A\subseteq\mathcal{X}$ is DS-shattered by $\mathcal{F}$ if there exists a finite $\mathcal{H}\subset\mathcal{F}$ such that for every $x\in A$ and $h\in\mathcal{H}$, there exists a $g\in\mathcal{H}$ such that $g(x)\neq h(x)$ and $g(z)=h(z)$ for all $z\in A\setminus\{x\}$. The DS dimension of $\mathcal{F}$, denoted $\emph{{DS}}(\mathcal{F})$, is the largest $d\in\mathbb{N}$ such that there exists a shattered set $A\subset\mathcal{X}$ with cardinality $d$. If there are arbitrarily large sets $A\subseteq\mathcal{X}$ that are shattered by $\mathcal{F}$, then we say that $\emph{{DS}}(\mathcal{F})=\infty.$
Definition 10 (Littlestone dimension [Littlestone, 1987, Daniely et al., 2011]).
Let $\mathcal{T}$ be a complete binary tree of depth $d$ whose internal nodes are labeled by a sequence $(\mathcal{T}_{0},\ldots,\mathcal{T}_{d-1})$ of node-labeling functions $\mathcal{T}_{t-1}:\{-1,1\}^{t-1}\to\mathcal{X}$. The tree $\mathcal{T}$ is shattered by $\mathcal{F}\subseteq\mathcal{X}^{\mathcal{X}}$ if there exists a sequence $(Y_{1},...,Y_{d})$ of edge-labeling functions $Y_{t}:\{-1,1\}^{t}\rightarrow\mathcal{X}$ such that for every path $\sigma=(\sigma_{1},...,\sigma_{d})\in\{-1,1\}^{d}$, there exists a function $f_{\sigma}\in\mathcal{F}$ such that for all $t\in[d]$, $f_{\sigma}(\mathcal{T}_{t-1}(\sigma_{<t}))=Y_{t}(\sigma_{\leq t})$ and $Y_{t}((\sigma_{<t},-1))\neq Y_{t}((\sigma_{<t},+1))$. The Littlestone dimension of $\mathcal{F}$, denoted $\emph{{L}}(\mathcal{F})$, is the maximal depth of a tree $\mathcal{T}$ that is shattered by $\mathcal{F}$. If there exists shattered trees of arbitrarily large depth, we say $\emph{{L}}(\mathcal{F})=\infty$.
@@ Appendix A Combinatorial dimensions
Proof.
(of lowerbound in (ii) of Theorem 3.1) Fix time horizon $T\in\mathbb{N}$ and scale $\gamma>0$. By definition of $\texttt{E}_{\gamma}(\mathcal{F})$, there exists a trajectory tree $\mathcal{T}$ of depth $\texttt{E}_{\gamma}(\mathcal{F})$ such that $\mathcal{B}(\mathcal{T})\geq\gamma\texttt{E}_{\gamma}(\mathcal{F})$. Let $\mathcal{T}^{\prime}$ denote the trajectory tree after removing the last $\max\{\texttt{E}_{\gamma}(\mathcal{F})-T,0\}$ levels from $\mathcal{T}$. Note that the depth of $\mathcal{T}^{\prime}$ is exactly $T$ and its branching factor $\mathcal{B}(\mathcal{T}^{\prime})\geq\gamma\texttt{E}_{\gamma}(\mathcal{F})-\max\{\texttt{E}_{\gamma}(\mathcal{F})-T,0\}.$ Thus, it must be the case that $\texttt{C}_{T}(\mathcal{F})\geq\gamma\texttt{E}_{\gamma}(\mathcal{F})-\max\{\texttt{E}_{\gamma}(\mathcal{F})-T,0\}$. Since $\gamma$ was chosen arbitrarily, this completes the proof. ∎
Proof.
(of upperbound in (ii) of Theorem 3.1). Fix a $\gamma>0$. Using Definition 6, we first establish two facts: (1) $\texttt{B}(\mathcal{T})<\gamma(\texttt{E}_{\gamma}(\mathcal{F})+1)$ for every shattered trajectory tree $\mathcal{T}$ of depth $d\leq\texttt{E}_{\gamma}(\mathcal{F})$ and (2) $\texttt{B}(\mathcal{T})<\gamma d$ for every shattered trajectory tree $\mathcal{T}$ of depth $d>\texttt{E}_{\gamma}(\mathcal{F})$. To see fact (1), if there exists a shattered trajectory tree $\mathcal{T}$ of depth $\leq\texttt{E}_{\gamma}(\mathcal{F}):=d$ such that $\texttt{B}(\mathcal{T})\geq\gamma(\texttt{E}_{\gamma}(\mathcal{F})+1)$, one can extend this tree to obtain a shattered trajectory tree $\mathcal{T}^{\prime}$ of depth $d+1$ with $\texttt{B}(\mathcal{T}^{\prime})\geq\gamma(d+1)$. This would imply that $\texttt{E}_{\gamma}(\mathcal{F})=d+1$, a contradiction. Fact (2) follows by a similar reasoning. If there exists a shattered trajectory tree $\mathcal{T}$ of depth $d>\texttt{E}_{\gamma}(\mathcal{F})$ such that $\texttt{B}(\mathcal{T})\geq\gamma d$, then $\texttt{E}_{\gamma}(\mathcal{F})\geq d$, leading to a contradiction.
Finally, to prove the upperbound in (ii) in Theorem 3.1, we can use facts (1) and (2) to write
$\texttt{C}_{T}(\mathcal{F})$ $=\texttt{C}_{T}(\mathcal{F})\mathbbm{1}\{T\leq\texttt{E}_{\gamma}(\mathcal{F})\}+\texttt{C}_{T}(\mathcal{F})\mathbbm{1}\{T>\texttt{E}_{\gamma}(\mathcal{F})\}$
$\leq\gamma(\texttt{E}_{\gamma}(\mathcal{F})+1)\mathbbm{1}\{T\leq\texttt{E}_{\gamma}(\mathcal{F})\}+\gamma T\mathbbm{1}\{T>\texttt{E}_{\gamma}(\mathcal{F})\}$
$\leq\gamma\max\{\texttt{E}_{\gamma}(\mathcal{F}),T\}+\gamma,$
Taking the infimum with respect to $\gamma>0$ completes the proof. ∎
Appendix C Proof of Theorem 3.2
The equivalence between statements (1) and (2) follows directly from the quantitative bounds in Theorem 3.1. Thus, we only focus on proving the equivalence between statements (2) and (3). The direction $(3)\implies(2)$ follows from upperbound (iv) in Theorem 3.1 and the fact that $\inf_{\gamma>0}\{\gamma\max\{\texttt{E}_{\gamma}(\mathcal{F}),T\}+\gamma\}=o(T)$. To see the later, observe that if $\texttt{E}_{\gamma}(\mathcal{F})<\infty$ for every $\gamma>0$, then
$\lim_{T\rightarrow\infty}\frac{\inf_{\gamma>0}\{\gamma\max\{\texttt{E}_{\gamma}(\mathcal{F}),T\}+\gamma\}}{T}\leq\inf_{\gamma>0}\Bigl{\{}\lim_{T\rightarrow\infty}\frac{\gamma\max\{\texttt{E}_{\gamma}(\mathcal{F}),T\}+\gamma}{T}\Bigl{\}}=\inf_{\gamma>0}\gamma=0$
We now prove $(2)\implies(3)$. Suppose (2) is true, then by definition of $o(\cdot)$, for every $\varepsilon>0$, there exists a $T_{\varepsilon}\in\mathbb{N}$ such that for all $T>T_{\varepsilon}$, we have that $\texttt{C}_{T}(\mathcal{F})<\varepsilon T$. If $\texttt{C}_{T}(\mathcal{F})<\varepsilon T$ for all $T>T_{\varepsilon}$, then $\texttt{B}(\mathcal{T})<\varepsilon T$ for every shattered trajectory tree $\mathcal{T}$ of depth $T>T_{\varepsilon}$. Thus, it must be the case that $\texttt{E}_{\varepsilon}(\mathcal{F})\leq T_{\varepsilon}<\infty$.
Appendix D Proof of Theorem 3.3
@@
We now show that $\texttt{C}_{T}(\mathcal{F})\leq\sup_{x_{0}\in\mathbb{Z}}|S\cap\{|x_{0}|,...,|x_{0}|+T-1\}|$. Let $\mathcal{T}$ be any trajectory tree of depth $T$ shattered by $\mathcal{F}$. It suffices to show that $\texttt{B}(\mathcal{T})\leq|S\cap\{|\mathcal{T}_{0}|,...,|\mathcal{T}_{0}|+T-1\}|$. Fix any path $\varepsilon\in\{-1,1\}^{T}$ down $\mathcal{T}$ and consider the sequence of states $\mathcal{T}_{0},...,\mathcal{T}_{T}(\varepsilon_{\leq T})$. Note that $\mathcal{T}_{t}((\varepsilon_{<t},-1))\neq\mathcal{T}_{t}((\varepsilon_{<t},+1))$ only if $|\mathcal{T}_{t-1}(\varepsilon_{<t})|\in S$. Thus,

### v2

@@ Appendix A Proofs for Realizable learnability
Starting with (a), let $t\in[T]$ such that $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})>0$ and $\hat{x}_{t}\neq x_{t}$. We need to show that $\operatorname{C}_{T-t}(V_{t+1},x_{t})<\operatorname{C}_{T-t+1}(V_{t},x_{t-1})$. Suppose for the sake of contradiction that $\operatorname{C}_{T-t}(V_{t+1},x_{t})\geq\operatorname{C}_{T-t+1}(V_{t},x_{t-1}):=d$. Then, by the prediction rule, we have that $\operatorname{C}_{T-t}(V^{\hat{x}_{t}}_{t},\hat{x}_{t})\geq\operatorname{C}_{T-t}(V^{x_{t}}_{t},x_{t})\geq d$. By definition of the instance-dependent Evolution complexity, we are guaranteed the existence of a trajectory tree $\mathcal{T}_{x_{t}}$ of depth $T-t$, rooted at $x_{t}$, shattered by $V_{t}^{x_{t}}$ with branching factor at least $d$ and a tree $\mathcal{T}_{\hat{x}_{t}}$ of depth $T-t$, rooted at $\hat{x}_{t}$, shattered by $V_{t}^{\hat{x}_{t}}$ with branching factor at least $d$. Consider a binary tree $\mathcal{T}$ whose root node $v$ is labeled by $x_{t-1}$ and left and right subtrees are $\mathcal{T}_{x_{t}}$ and $\mathcal{T}_{\hat{x}_{t}}$ respectively. Then, observe that $\mathcal{T}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $d+1$ (because $x_{t}\neq\hat{x}_{t}$). This contradicts our assumption that $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})=d$. Thus, we must have $\operatorname{C}_{T-t}(V_{t+1},x_{t})<\operatorname{C}_{T-t+1}(V_{t},x_{t-1})$.
@@
Moving to (b), let $t\in[T]$ be such that $\hat{x}_{t}=x_{t}$. Then, we need to show that $d:=\operatorname{C}_{T-t}(V_{t+1},x_{t})\leq\operatorname{C}_{T-t+1}(V_{t},x_{t-1}).$ If $d=0$, the inequality holds trivially. Thus, assume $d>0$. Then, by definition, there exists a trajectory tree $\mathcal{T}$ of depth $T-t$, rooted at $x_{t}$, shattered by $V_{t+1}$ with branching factor at least $d$. Consider a binary tree $\tilde{\mathcal{T}}$ whose root node $v$ is labeled by $x_{t-1}$ and left and right subtrees are both $\mathcal{T}$. Then since $V_{t+1}=V_{t}^{x_{t}}\subseteq V_{t}$, we have that $\tilde{\mathcal{T}}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $d$. Thus, we must have $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})\geq d$.
@@
Finally for (c), let $t\in[T]$ be such that $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})=0$. Using the same logic as (b), $\operatorname{C}_{T-t}(V_{t+1},x_{t})=0$. Thus, to prove that (3) holds, it suffices to show that $x_{t}=\hat{x}_{t}$. To do so, we will show that the projection size |$|\{f(x_{t-1}):f\in V_{t}\}|=1$. Thus, Algorithm 1 does not make a mistake. Suppose for the sake of contradiction that |$|\{f(x_{t-1}):f\in V_{t}\}|\geq 2$. Then, there exists two functions $f_{1},f_{2}\in V_{t}$ such that $f_{1}(x_{t-1})\neq f_{2}(x_{t-1})$. Consider a binary tree $\mathcal{T}$ of depth $T-t+1$ where the root node is labeled by $x_{t-1}$, and its left and right child by $f_{1}(x_{t-1})$ and $f_{2}(x_{t-1})$ respectively. Label all remaining internal nodes in the left subtree of the root node such that every path is shattered by $f_{1}$. Likewise, for the right subtree of the root node, label every internal node such that every path is shattered by $f_{2}$. Following this procedure, $\mathcal{T}$ is a trajectory tree of depth $T-t+1$, rooted at $x_{t-1}$, shattered by $V_{t}$ with branching factor at least $1$ (because $f_{1}(x_{t-1})\neq f_{2}(x_{t-1})$). Thus, $\operatorname{C}_{T-t+1}(V_{t},x_{t-1})\geq 1$, which is a contradiction.
@@
We now show that $\operatorname{C}_{T}(\mathcal{F})\leq\sup_{x_{0}\in\mathbb{Z}}|S\cap\{|x_{0}|,...,|x_{0}|+T-1\}|$. Let $\mathcal{T}$ be any trajectory tree of depth $T$ shattered by $\mathcal{F}$. It suffices to show that $\operatorname{B}(\mathcal{T})\leq|S\cap\{|\mathcal{T}_{0}|,...,|\mathcal{T}_{0}|+T-1\}|$. Fix any path $\varepsilon\in\{-1,1\}^{T}$ down $\mathcal{T}$ and consider the sequence of states $\mathcal{T}_{0},...,\mathcal{T}_{T}(\varepsilon_{\leq T})$. Note that $\mathcal{T}_{t}((\varepsilon_{<t},-1))\neq\mathcal{T}_{t}((\varepsilon_{<t},+1))$ only if $|\mathcal{T}_{t-1}(\varepsilon_{<t})|\in S$. Thus,